let the harmonic mean and geometric mean of 2 +ve number be in the ratio 4:5 then the two numbers are in what ratio?

Grade:

1 Answers

Sher Mohammad IIT Delhi
9 years ago
harmonic mean 2ab/(a+b)
geometric mean sqrt(ab)
so 2ab/(a+b)*sqrt(ab)=4/5
ab/(a+b)2=4/25
(a/b)/(a/b +1)2=4/25
solving the quadratic equation we will get a/b=1/4 or 4 that means no. are in ratio 1:4

Sher Mohammad
faculty askiitias